# Learnability is a Compact Property

### v1

@@ Abstract
Recent work on learning has yielded a striking result: the learnability of various problems can be undecidable, or independent of the standard ZFC axioms of set theory. Furthermore, the learnability of such problems can fail to be a property of finite character: informally, it cannot be detected by examining finite projections of the problem.
On the other hand, learning theory abounds with notions of dimension that characterize learning and consider only finite restrictions of the problem, i.e., are properties of finite character. How can these results be reconciled? More precisely, which classes of learning problems are vulnerable to logical undecidability, and which are within the grasp of finite characterizations?
We demonstrate that the difficulty of supervised learning with metric losses admits a tight finite characterization. In particular, we prove that the sample complexity of learning a class $\mathcal{H}$ can be detected by examining its finite projections. For realizable and agnostic learning with respect to a wide class of proper loss functions, we demonstrate an exact compactness result: a class is learnable with a given sample complexity precisely when the same is true of all its finite projections. For realizable learning with improper loss functions, we show that exact compactness of sample complexity can fail, and provide matching upper and lower bounds of a factor of 2 on the extent to which such sample complexities can differ. We conjecture that larger gaps are possible for the agnostic case.
At the heart of our technical work is a compactness result concerning assignments of variables that maintain a class of functions below a target value, which generalizes Hall’s classic matching theorem and may be of independent interest.
@@ 1 Introduction
The vulnerabilities of mathematics were perhaps first demonstrated by Gödel’s revolutionary incompleteness theorems. Roughly speaking, they demonstrated that any reasonably expressive formal system suffers from the existence of statements which can be neither proven nor disproven [22]. Such statements are referred to as being undecidable, and for the standard axioms of ZFC, they have been shown to reach across varied fields of mathematics, including set theory, topology, algebra, and beyond [17, 34, 38, 33]. [^1] We emphasize that we use the term “undecidable” in the logical sense rather than the computational sense, i.e., to mean independent of the axioms of ZFC set theory. We do not discuss computational undecidability in this paper.
It was recently demonstrated that problems in learning theory can face a similar fate. Ben-David et al. [10] constructed a hypothesis class $\mathcal{H}$ in the setting of EMX learning, a variant of binary classification, whose learnability is undecidable provided that learners are obligated to emit hypotheses in $\mathcal{H}$. In the years since, related undecidability results have also been established in reinforcement learning and online learning [26, 16].
An important consequence of the work of Ben-David et al. [10] is that the learnability of EMX problems is not a property of finite character, meaning it cannot be detected by examining finite projections of a problem (i.e., finite subsets of the domain and of the hypothesis class). Learning theory, however, is rich with characterizations of learnability that are indeed properties of finite character: the VC dimension for binary classification [13], the DS dimension for multiclass classification [14], and the fat-shattering dimension for real-valued functions [9, 5], to name a few. Formally, for each such dimension $D(\mathcal{H})$, one can certify that a class $\mathcal{H}$ has $D(\mathcal{H})\geq d$ using only a finite collection of domain points and a finite subset of $\mathcal{H}$. (E.g., for the VC dimension, a collection of $d$ domain points which are shattered by $\mathcal{H}$ and $2^{d}$ functions in $\mathcal{H}$ which shatter them.)
How, then, are such results to be reconciled with the undecidability of EMX learning? More precisely, what are the particular features of EMX learning which prevent it from enjoying a characterization of learnability in the style of classification and regression? (Or, conversely, what special properties of classification and regression grant it the fortune of admitting finite characterizations?) Existing research on undecidability has largely taken an individualized approach toward understanding particular classes of problems, shedding comparatively less light on the previous questions. The purpose of this paper, in contrast, is to address the following fundamental question:
For which classes of supervised learning problems can the difficulty of learning a hypothesis class $\mathcal{H}$ be detected by examining its finite projections? Why?
Notably, we take the perspective of considering not only learnability but also the sample complexity of learning, i.e., the volume of data required to attain favorable performance.
As our cornerstone result, we demonstrate that for the case of supervised learning with a proper label space — including $\mathbb{R}^{n}$ and its closed subsets, under any norm — a class $\mathcal{H}$ can be learned with sample complexity $m$ precisely when the same is true of all its finite projections. Notably, we demonstrate the result not only for learning in the realizable and agnostic cases, but also for distribution-family learning with respect to a wide class of well-behaved distribution classes, including those used in EMX learning and the learning of partial concept classes [6]. For learning over arbitrary label spaces, we demonstrate that this property fails. In the realizable case, we provide matching upper and lower bounds of a factor of 2 on the extent to which such sample complexities can differ. Our lower bound transfers directly to learning in the agnostic case, for which we conjecture that larger gaps in sample complexity are possible.
Notably, our results show that supervised learning with metric losses is immune to the pathology of EMX learning: it cannot be that a class $\mathcal{H}$ is undecidable to learn (or even difficult to learn), provided that all its finite projections are easily learned. Informally, there cannot occur a “phase change” from learning finite classes to infinite ones (or between infinite classes of varying cardinalities, as occurs in [10]; see Figure 1). This demonstrates that the result of [10] relied crucially upon the requirement that learners only emit hypotheses in $\mathcal{H}$, which we do not adopt in this work. That is, it demonstrates a structural difference between learnability with respect to proper learners, which are constrained to emit hypotheses in $\mathcal{H}$, and improper learners, which are not. (We note that we will shortly be overloading the term “proper”, as we will be introducing the notions of proper metric spaces and proper functions between metric spaces. We trust the reader to avoid confusion.)
Our main result involves a generalization of the classic marriage theorems for bipartite graphs, which might render it of independent mathematical interest. The original marriage theorem, due to Philip Hall [23], articulates a necessary and sufficient condition for the existence of a perfect matching of one side of a finite bipartite graph into the other. Subsequently, Marshall Hall [24] extended the same characterization of finite character to infinite graphs of arbitrary cardinality (even uncountable), provided the side to be matched has finite degrees — the characterization being false otherwise, as can be seen by a simple countable example. The proof of M. Hall features an involved analysis of the lattice of “blocking sets”, and invokes the axiom of choice through Zorn’s lemma. Simpler proofs of the same result have since been discovered: a topological proof by [25] which invokes the axiom of choice through Tychonoff’s theorem, and an algebraic proof by [35] which also uses Zorn’s lemma. At the heart of our main result is a characterization of finite character for a variable-assignment problem which generalizes both supervised learning and bipartite matching: one side of a bipartite graph indexes infinitely many variables, the other indexes infinitely many functions that depend on finitely many variables each, and the goal is to assign all the variables in a manner that maintains all functions below a target value. Our proof draws inspiration from all three of the aforementioned proofs of M. Hall’s theorem, and goes through Zorn’s lemma.
[…]
Figure 1: Learnability of a hypothesis class $\mathcal{H}$, provided that all its finite projections are easily learned. Green regions denote cardinalities of learnable classes and red regions of nonlearnable classes. In “classical learning” — e.g., binary classification, multiclass classification with arbitrary label sets, and regression — such a class $\mathcal{H}$ is guaranteed to be learnable regardless of its cardinality, $|\mathcal{H}|$. In EMX learning, Ben-David et al. [10] exhibit a class whose finite projections are trivial to learn yet which is learnable precisely when $|\mathcal{H}|<\aleph_{\omega}$, i.e., when there exists $k\in\mathbb{N}$ for which $\mathcal{H}\leq\aleph_{k}$ (from which it follows that the learnability of a continuum-sized class is undecidable). Informally, EMX learning exhibits a “phase change” in which sufficiently large infinite classes $\mathcal{H}$ may suddenly fail to be learnable. For classical learning, in contrast, the local structure of $\mathcal{H}$ (i.e., its family of finite projections) reveals complete information concerning its learnability: there are no sudden “phase changes.”
1.1 Related Work
Ben-David et al. [10] were the first to demonstrate that learning problems can be undecidable, by way of a problem in EMX (Estimating the Maximum) learning. Subsequently, Hanneke and Yang [26] demonstrated that the same can occur for problems in reinforcement learning, while Caro [16] proved the existence of undecidable problems in binary classification and online learning. Notably, Caro [16] considers the case in which the learnability of $\mathcal{H}$ must be decided given only a form of black-box query access to $\mathcal{H}$. That is, the elements of $\mathcal{H}$ are indexed, possibly redundantly, by values $i\in I$ and one is permitted to evaluate $f_{i}$ at the datapoint $x$, for $(i,x)$ of one’s choosing. In this setting, Caro [16] demonstrates that it is undecidable merely to distinguish whether $|\mathcal{H}|=1$ or $|\mathcal{H}|=\infty$! Our work, in constrast, assumes complete knowledge of $\mathcal{H}$ (or its projections) as sets.
Hart [27] demonstrated that the result of Ben-David et al. [10] is sensitive to the requirement that learners be measurable: in particular, such an undecidability result could not hold when learners are required to have finite representations, e.g., to be computable. Thus Agarwal et al. [4] initiated a study of computable PAC (CPAC) learning, in which learners are required to be algorithmic. Subsequent work investigating the relationship between PAC and CPAC learning includes Ackerman et al. [1], Delle Rose et al. [20], Hasrati and Ben-David [28], and Sterkenburg [39]. Lathrop [30] previously studied the learnability of noncomputable problems, a distinct setting. The computational (un)decidability of learning — e.g., the ability of a Turing machine to decide whether a class is learnable — has also received attention [36, 12, 15], but bears little direct relation to our work.
David et al. [19] established a loose compactness result for the special case of multiclass classification, by way of a sample compression argument. A similar result is implied by the characterization of multiclass learnability through the DS dimension [18, 14] and subsumed by our exact compactness result in Corollary 3.13 (see also Theorem 3.15). We note that our results extend well beyond multiclass classification, across all of supervised learning with metric losses.
Perhaps most related to the present work is Attias et al. [8], which introduces the $\gamma$-OIG dimension and demonstrates that it characterizes learnability for supervised learning problems with pseudometric losses. Notably, this is the first general dimension characterizing learnability across essentially the entirety of supervised learning. The $\gamma$-OIG dimension itself serves to rebut the undecidability result of Ben-David et al. [10], but we note that it has not been shown to tightly characterize the sample complexity of learning, and furthermore that it is analyzed only for realizable learning. Our work, in contrast, provides a nearly complete characterization regarding the sample complexity of learning for both the realizable and agnostic cases. In fact, in Section 3.4 we extend our results to certain cases of distribution-family learning, which permits us to shed light on learning paradigms such as partial concept classes and EMX learning itself, the source of all the excitement.
2 Preliminaries
2.1 Notation
For a natural number $n\in\mathbb{N}$, $[n]$ denotes the set $\{1,\ldots,n\}$. For a predicate $P$, $[P]$ denotes the Iverson bracket of $P$, i.e.,
$[P]=\begin{cases}1&\text{$P$ is true},\\ 0&\text{$P$ is false.}\end{cases}$
When $Z$ is a set, $Z^{<\omega}$ denotes the set of all finite sequences in $Z$, i.e., $Z^{<\omega}=\bigcup_{i=1}^{\infty}Z^{i}$. For a tuple $S=(z_{1},\ldots,z_{n})$, we use $S_{-i}$ to denote $S$ with its $i$th entry removed, i.e.,
$S_{-i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n}).$
2.2 PAC Learning
Let us recall the standard toolkit of supervised learning. A learning problem is determined by a domain $\mathcal{X}$, label space $\mathcal{Y}$, and hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$. The elements of $\mathcal{H}$ are functions $\mathcal{X}\to\mathcal{Y}$, and are referred to as hypotheses or predictors. Learning also requires a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$, which usually endows $\mathcal{Y}$ with the structure of a metric space. Throughout the paper, we permit $\mathcal{X}$ to be arbitrary. A labeled datapoint is a pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and an unlabeled datapoint is an element $x\in\mathcal{X}$. We may refer to labeled and unlabeled datapoints simply as datapoints when clear from context.
A training set, or tranining sample, is a tuple of labeled datapoints $S\in(\mathcal{X}\times\mathcal{Y})^{<\omega}$. A learner is a function from training sets to hypotheses, i.e., $A:(\mathcal{X}\times\mathcal{Y})^{<\omega}\to\mathcal{Y}^{\mathcal{X}}$. The purpose of a learner $A$ is to use the information contained in $S$ to produce a function $A(S):\mathcal{X}\to\mathcal{Y}$ that faithfully captures the underlying relationship between datapoints and their labels. Such an underlying relationship is formalized as a probability measure $D$ over $\mathcal{X}\times\mathcal{Y}$, which we may refer to as a ground truth. The true error, or simply error, incurred by a predictor $h$ with respect to $D$ is the average loss it suffers over datapoints drawn from $D$, i.e.,
$L_{D}(h)=\operatorname*{\mathbb{E}}_{(x,y)\sim D}\ell(h(x),y).$
The success or failure of a learner is formalized by way of Valiant’s celebrated Probability Approximately Correct (PAC) learning paradigm [40].
Definition 2.1.
$L_{D}(A(S))\leq\inf_{\mathcal{H}}L_{D}(h)+\epsilon.$
When $\mathbb{D}$ consists of all probability measures over $\mathcal{X}\times\mathcal{Y}$, then $A$ is an agnostic PAC learner for $\mathcal{H}$. When $\mathbb{D}$ consists of all probability measure $D$ such that $L_{D}(h)=0$ for some $h\in\mathcal{H}$, then $A$ is a realizable PAC learner for $\mathcal{H}$.
Definition 2.2.
The sample complexity of a learner $A$ with respect to a hypothesis class $\mathcal{H}$, $m_{\operatorname{PAC},A}:(0,1)^{2}\to\mathbb{N}$, is the minimal sample function it attains as a learner for $\mathcal{H}$. The sample complexity of a class $\mathcal{H}$ is the pointwise minimal sample complexity attained by any of its learners,
$m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\delta)=\min_{A}m_{\operatorname{PAC},A}(\epsilon,\delta).$
We remark briefly that we adopt a rather hands-off approach regarding measure-theoretic details, and note that it suffices to adopt the mild measurability assumptions of, e.g., [37] in order to ensure that true error and PAC learnability be well-defined. See e.g. [13], [32, Appendix C], or Dudley et al. [21, Section 10.3.1] for a more precise measure-theoretic treatment of function classes and learnability.
2.3 Transductive Learning
The paradigm of transductive learning stands in contrast to Valiant’s classic PAC learning, by demanding that learners attain favorable performance on each sample rather than each distribution. Notably, it dates to at least the pioneering work of Haussler et al. [29] and has witnessed renewed interest in recent years owing to the mathematical insights provided by one-inclusion graphs for capturing transductive learning (see, e.g., [18, 6, 14, 3, 2, 7, 8]).
Crucially, despite the seeming difficulty of transductive learning relative to PAC learning, they amount to essentially equivalent benchmarks for assessing learners. This justifies our emphasis on transductive learnability and sample complexity in the forthcoming sections.
Definition 2.3.
The setting of transductive learning is defined by the following sequence of steps:
@@ 2.3 Transductive Learning
    An adversary selects $S=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ and a hypothesis $h\in\mathcal{H}$.
@@
    and their labels under $h$ are displayed to the learner. That is, the learner receives $\big{(}x_{j},h(x_{j})\big{)}_{x_{j}\in S_{-i}}$.

### v2

# Transductive Sample Complexities Are Compact

@@ Abstract
We demonstrate a compactness result holding broadly across supervised learning with a general class of loss functions: Any hypothesis class $\mathcal{H}$ is learnable with transductive sample complexity $m$ precisely when all of its finite projections are learnable with sample complexity $m$. We prove that this exact form of compactness holds for realizable and agnostic learning with respect to any proper metric loss function (e.g., any norm on $\mathbb{R}^{d}$) and any continuous loss on a compact space (e.g., cross-entropy, squared loss). For realizable learning with improper metric losses, we show that exact compactness of sample complexity can fail, and provide matching upper and lower bounds of a factor of 2 on the extent to which such sample complexities can differ. We conjecture that larger gaps are possible for the agnostic case. Furthermore, invoking the equivalence between sample complexities in the PAC and transductive models (up to lower order factors, in the realizable case) permits us to directly port our results to the PAC model, revealing an almost-exact form of compactness holding broadly in PAC learning.
@@ 1 Introduction
Compactness results in mathematics describe the behavior by which, roughly speaking, an infinite system can be entirely understood by inspecting its finite subsystems: An infinite graph is $k$-colorable precisely when its finite subgraphs are all $k$-colorable (De Bruijn and Erdös, 1951), an infinite collection of compact sets in $\mathbb{R}^{d}$ has non-empty intersection precisely when the same is true of its finite subcollections, etc. In each case, compactness reveals a profound and striking structure, by which local understanding of a problem immediately yields global understanding.
We demonstrate that supervised learning in the transductive model enjoys such structure. First, let us briefly review the transductive model, a close relative of the PAC model. In the realizable setting with a class of hypotheses $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, it is defined by the following sequence of steps:
@@
    An adversary selects unlabeled data $S=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ and a hypothesis $h\in\mathcal{H}$.
@@
    and their labels under $h$ are displayed to the learner.
@@
The expected error incurred by the learner over the uniformly random choice of $x_{i}$ is its transductive error on this learning instance, from which one can easily define the transductive sample complexity of a learner and of a hypothesis class.
Notably, transductive learning, originally introduced by Vapnik and Chervonenkis (1974) and Vapnik (1982), is a fundamental approach to learning with deep theoretical connections to the PAC model. We study the transductive model as employed by the pioneering work of Haussler et al. (1994), who introduced the celebrated one-inclusion graph (OIG) to study transduction and used it to derive improved error bounds for VC classes. More recently, transductive learning and OIGs have been used to (among other work) establish the first characterizations of learnability for multiclass classification and realizable regression (Brukhim et al., 2022; Attias et al., 2023), to prove optimal PAC bounds across several learning settings (Aden-Ali et al., 2023b), and to understand regularization in multiclass learning (Asilis et al., 2024). (See also Daniely and Shalev-Shwartz (2014); Alon et al. (2022); Montasser et al. (2022); Aden-Ali et al. (2023a).) The transductive model also naturally generalizes to the agnostic setting, much like PAC learning, as articulated by Asilis et al. (2024).
1.1 Contributions
Our results involve comparing a hypothesis class $\mathcal{H}$ to its “finite projections.” Formally, for a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and finite collection of unlabeled data $S\subseteq\mathcal{X}$, we refer to the finite subsets of $\mathcal{H}|_{S}$ as finite projections of $\mathcal{H}$. Note that $\mathcal{H}$ is being “made finite” at two levels: first by restricting its functions to a finite region $S\subseteq\mathcal{X}$ of the domain, and second by passing to a finite subset of $\mathcal{H}|_{S}$.
As our cornerstone result, we demonstrate in Theorem 3.6 that for the case of supervised learning with a large class of proper [^1] We warn that we will shortly be overloading the term “proper”, as we discuss proper metric spaces and proper functions between metric spaces. We also note that our notion of properness is unrelated to losses which incentivize predicting the true probability, from e.g. Blasiok et al. (2023). (And unrelated to proper vs. improper learners; we consider improper learners throughout the paper, which can emit predictors outside the class $\mathcal{H}$.) metric loss functions (including any norm on $\mathbb{R}^{d}$ or its closed subsets; see Definition 3.2) a class $\mathcal{H}$ can be learned with transductive sample complexity $m$ precisely when the same is true of all its finite projections. In fact, in Theorem 3.7 we extend our results to arbitrary continuous losses on compact metric spaces, e.g., cross-entropy loss on finite-dimensional probability spaces and squared $\ell_{2}$ loss on compact subsets of $\mathbb{R}^{d}$. For learning over arbitrary label spaces, we demonstrate in Theorems 3.8 and 3.9 that compactness fails: for realizable learning with metric losses, we provide matching upper and lower bounds of a factor of 2 on the extent to which such transductive sample complexities can differ. Our lower bound transfers directly to transductive learning in the agnostic case, for which we conjecture that larger gaps in sample complexity are possible.
We stress that our compactness results are exact in the transductive model, avoiding dilution by asymptotics or even by constants. In addition, there is a growing body of work relating sample complexities in the transductive and PAC models, by which our results directly transfer in a black-box manner (Asilis et al., 2024; Aden-Ali et al., 2023b; Dughmi et al., 2024). Notably, for realizable learning with any bounded loss, PAC sample complexities differ from their transductive counterparts by at most a logarithmic factor in $\delta$, the confidence parameter. Combined with our results, this reveals an almost-exact form of compactness for realizable PAC learning, as we describe in Section 3.4.222Note too that any future improvements to the connections between the PAC and transductive models, whether in the realizable or agnostic settings, will be automatically inherited by our results in a black-box manner.
Our results hold for improper learners, i.e., learners that are permitted to emit a predictor outside the underlying class $\mathcal{H}$. Curiously, compactness of sample complexity can be seen to fail strongly when one requires that learners be proper, using the work of Ben-David et al. (2019). This demonstrates a structural difference between proper and improper learning; see Section 3.5 for further detail.
Our compactness results are underpinned by a generalization of the classic marriage theorems for bipartite graphs which may be of independent mathematical interest. The original marriage theorem, due to Philip Hall (Hall, 1935), articulates a necessary and sufficient condition for the existence of a perfect matching from one side of a finite bipartite graph to the other. Subsequently, Marshall Hall (Hall Jr, 1948) extended the same characterization, referencing only finite subgraphs, to infinite graphs of arbitrary cardinality, provided the side to be matched has finite degrees — the characterization being false otherwise, as can be seen by a simple countable example. This characterization therefore serves as a compactness result for matching on such infinite graphs. The proof of M. Hall features an involved analysis of the lattice of “blocking sets”, and invokes the axiom of choice through Zorn’s lemma. Simpler proofs have since been discovered: a topological proof by Halmos and Vaughan (1950) which invokes the axiom of choice through Tychonoff’s theorem, and an algebraic proof by Rado (1967) which also uses Zorn’s lemma. At the heart of our paper is a compactness result (Theorem 3.3) for a variable-assignment problem which generalizes both supervised learning and bipartite matching: one side of a bipartite graph indexes infinitely many variables, the other indexes infinitely many functions that depend on finitely many variables each, and the goal is to assign all the variables in a manner that maintains all functions below a target value. Our proof draws inspiration from all three of the aforementioned proofs of M. Hall’s theorem, and goes through Zorn’s lemma.
1.2 Related Work
The transductive approach to learning dates to the work of Vapnik and Chervonenkis (1974) and Vapnik (1982), and has inspired a breadth of recent advances across regression, classification, and various other learning regimes; see our introduction for a brief overview. Regarding transductive sample complexities, Hanneke et al. (2023) recently demonstrated a trichotomy result for optimal transductive error rates in the online setting of Ben-David et al. (1997). In contrast, we focus on the classical (batch) setting, as described in Section 2.2.
Perhaps most related to the present work is Attias et al. (2023), which introduces the $\gamma$-OIG dimension and demonstrates that it characterizes learnability for supervised learning problems with pseudometric losses. Notably, this is the first general dimension characterizing learnability across essentially the entirety of supervised learning. The $\gamma$-OIG dimension itself establishes a qualitative form of compactness — as it is defined using only the finite projections of a class — but we note that it has not been shown to tightly characterize the sample complexity of learning. Furthermore, it is analyzed only for realizable learning, which is in general not equivalent to agnostic learning (e.g., for regression). Our work, in contrast, establishes exact compactness for the sample complexity of transductive learning for both the realizable and agnostic settings, with respect to a general class of loss functions. Moreover, in Section 3.5 we extend our results to certain cases of distribution-family learning, including realizable learning of partial concept classes.
2 Preliminaries
2.1 Notation
For a natural number $n\in\mathbb{N}$, $[n]$ denotes the set $\{1,\ldots,n\}$. For a predicate $P$, $[P]$ denotes the Iverson bracket of $P$, i.e., $[P]=1$ when $P$ is true and 0 otherwise. When $Z$ is a set, $Z^{<\omega}$ denotes the set of all finite sequences in $Z$, i.e., $Z^{<\omega}=\bigcup_{i=1}^{\infty}Z^{i}$. For a tuple $S=(z_{1},\ldots,z_{n})$, we use $S_{-i}$ to denote $S$ with its $i$th entry removed, i.e., $S_{-i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})$.
2.2 Transductive Learning
Let us recall the standard toolkit of supervised learning. A learning problem is determined by a domain $\mathcal{X}$, label space $\mathcal{Y}$, and hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$. The elements of $\mathcal{H}$ are functions $\mathcal{X}\to\mathcal{Y}$; such functions are referred to as hypotheses or predictors. Learning also requires a loss function $\ell$ (or $d$) from $\mathcal{Y}\times\mathcal{Y}$ to $\mathbb{R}_{\geq 0}$, which often endows $\mathcal{Y}$ with the structure of a metric space. Throughout the paper, we permit $\mathcal{X}$ to be arbitrary. A labeled datapoint is a pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and an unlabeled datapoint is an element $x\in\mathcal{X}$. A training set, or training sample, is a tuple of labeled datapoints $S\in(\mathcal{X}\times\mathcal{Y})^{<\omega}$. A learner is a function from training sets to predictors, i.e., $A:(\mathcal{X}\times\mathcal{Y})^{<\omega}\to\mathcal{Y}^{\mathcal{X}}$.
Definition 2.1.
Realizable transductive learning is defined as follows: An adversary selects $S=(x_{i})_{i\in[n]}\in\mathcal{X}^{n}$ and a hypothesis $h\in\mathcal{H}$. The unlabeled datapoints $S$ are displayed to the learner. Then one datapoint $x_{i}$ is selected uniformly at random from $S$, and the remaining datapoints and their labels under $h$ are displayed to the learner. Lastly, the learner is prompted to predict the label of $x_{i}$, i.e., $h(x_{i})$.